OpenStudy (grant330sims):

Find the angle between the 2 Vectors: u=<2,-3> v=<1,-2>

OpenStudy (anonymous):

Now we should use this formula \[\alpha=arcos\left(\frac{\overrightarrow{u}\cdot\overrightarrow{v}}{||\overrightarrow{u}||*||\overrightarrow{v}||}\right)\] \[\overrightarrow{u}\cdot\overrightarrow{v}=2*1+(-3)*(-2)\] \[\overrightarrow{u}\cdot\overrightarrow{v}=8\] \[||\overrightarrow{u}||=\sqrt{2^2+(-3)^2}=\sqrt{13}\] \[||\overrightarrow{v}||=\sqrt{1^2+(-2)^2}=\sqrt{5}\] replace this values \[\alpha=arcos\left(\frac{8}{\sqrt{13}*\sqrt{5}}\right)\] \[\boxed{\boxed{\alpha\approx0.78\approx44.78^o}}\]

OpenStudy (grant330sims):

#ComingInClutch

OpenStudy (anonymous):

;D